### v1

@@ 4 Conclusions and Remarks
The search for life in our Universe is a vast arena of study. In this work, we propose the idea of searching for extraterrestrial artificial intelligence (SET-AI) as an advanced civilization using the proposal that ET-AI uses interstellar Dyson sphere-like mega-structures to harvest energy. For the measure of civilization advancement, we introduce the space exploration length. We propose a relation between this length and the Kardashev scale. One step further, we argue that primordial black holes as plausible candidates of DM are distributed evenly in the DM halo. They are more abundant and efficient energy sources than the main sequence stars for reaching and harvesting energy beyond the star-planet system. We investigate the percentage of Kardashev scale achievement as a function of the fraction of PBH as DM and the mass of PBH shown in figure 1 and 2.
We explore the prospects of detecting ET-AI. On a galactic scale, one can search for Infrared and submillimeter excess in the interstellar medium plus teh feature of a multi-blackbody spectrum. ALMA is a suitable telescope for searching these structures, even for direct imaging up to scales of $\sim 5.4\rm{kpc}$. These investigations apply to the galaxies beyond our Milky Way, and we can consider the modified Drake equation. Beyond the Milky Way with the ALMA telescope, one can search for a Dyson sphere-like structure up to $2\rm{Mpc}$ and for higher redshifts, one can look at the excess Infrared radiation in a plot versus the UV-continuum slope of a galaxy. In future studies in this direction, we will focus on the dependency of the accretion efficiency of PBH on their mass, redshift and environment. The galaxy properties can also affect the results of the achieved Kardashev parameter. The observational hints of this idea need further and extensive investigation. In conclusion, it seems that it is worth paying attention to finding ET-AI when we are looking for life in the cosmos.
@@ ACKNOWLEDGMENTS
SB is partially supported by the Abdus Salam International Center for Theoretical Physics (ICTP) under the regular associateship scheme. SB is partially supported by the Sharif University of Technology Office of Vice President for Research under Grant No. G4010204.

### v2

@@ 4 Conclusions and Remarks
The search for life in our Universe is a vast arena of study. In this work, we propose the idea of SET-AI as an advanced civilization using the proposal that ET-AI uses interstellar Dyson sphere–like megastructures to harvest energy. For the measure of civilization advancement, we introduce the space exploration length, or SED. We propose a relation between this length and the Kardashev scale. One step further, we argue that PBHs as plausible candidates of DM, are distributed evenly in the DM halo. They are more abundant and efficient energy sources than the main–sequence stars for reaching and harvesting energy beyond the star–planet system. We investigate the percentage of Kardashev scale achievement as a function of the fraction of PBHs as DM and the masses of PBHs, as shown in Figures 1 and 2.
We explore the prospects of detecting ET-AI. On a galactic scales, one can search for infrared and submillimeter excess in the interstellar medium plus the feature of a multiblackbody spectrum. ALMA is a suitable telescope for searching for these structures, even for direct imaging up to scales of $\sim 5.4\rm{kpc}$. These investigations apply to the galaxies beyond our Milky Way, and we can consider the modified Drake equation. Beyond the Milky Way with the ALMA telescope, one can search for a Dyson sphere–like structure up to $2\rm{Mpc}$ and for higher redshifts, one can look at the excess infrared radiation in a plot versus the UV continuum slope of a galaxy. In future studies in this direction, we will focus on the dependency of the accretion efficiency of PBHs on their mass, redshift, and environment. The galaxy properties can also affect the results of the achieved Kardashev parameter. The observational hints of this idea need further and extensive investigation. In conclusion, it seems that it is worth paying attention to finding ET-AI when we are looking for life in the cosmos.
@@ ACKNOWLEDGMENTS
S.B. is partially supported by the Abdus Salam International Center for Theoretical Physics (ICTP) under the regular associateship scheme. S.B. is partially supported by the Sharif University of Technology Office of Vice President for Research under grant No. G4010204.